# Enhanced Generative Data Augmentation for Semantic Segmentation via Stronger Guidance

### v3

@@ 1 INTRODUCTION
  • Visual Prior Blending: This method combines the visual priors from the original image and the segmentation masks. It aims to balance the information between the original image and the labeled objects, ensuring that the generated objects maintain the structure defined by the segmentation labels.
@@ 3 METHOD
Our proposed image data augmentation pipeline is shown in Figure 1, which consists of three main components: (1) Text prompt construction, (2) Visual Prior Blending, and (3) Controllable Diffusion Generation. Class-Prompt Appending append “class prompt” including the classes visible in the image with “caption” generated from the Image Captioning model. Visual Prior Blending method combines the visual pre-information of the real image and the segmentation map. The results of the above two methods are fed into the Controllable Steady Diffusion model to generate the synthetic image. In addition, to generate synthetic data from a given dataset, we use class balancing algorithm to generate data with even distribution among classes. Next, we provide a detailed description of each component.
@@ 3.2.2 Visual Prior Blending
In general, methods such as Line-Art Edge, Canny Edge, or HED soft edge all suffer from the limitation that the labeled classes in the image may be blurred or small in size, leading to inaccuracies in describing the structure of the labeled classes within the conditional image. Figure 4(a) shows an image produced using Line Art Detection. However, the edge results in this case are missing some details of the person, and the TV monitor is almost absent. The red shapes indicate the missing details in the image. This loss of information also occurs when using HED or Canny Edge. These weaknesses result in mislabeling in the synthetic image compared to the original image. We observed that although the segmentation labels of real images cannot fully describe an image’s content, they provide accurate information about the labeled classes. Based on this observation, we propose combining the real image’s prior visualization with the labels before feeding them into the controlled image generation model. The combination of the prior visual image $\mathcal{I}_{i}$ ($\mathcal{V}^{\mathcal{I}}_{i}$) and prior visual segmentation label $\mathcal{S}_{i}$ ($\mathcal{V}^{\mathcal{S}}_{i}$) ensures that the generated image has a clear layout and well preserves structures the class labeled information. Our proposed combines $\mathcal{V}^{\mathcal{I}}_{i}$ and $\mathcal{V}^{\mathcal{S}}_{i}$ by a weighted sum:
@@
With $\omega_{1}$, $\omega_{2}$ being the trade-off scales when combining $\mathcal{V}^{\mathcal{I}}_{i}$ and $\mathcal{V}^{\mathcal{S}}_{i}$. This combination results in a prior visual that is clear in content and complete information about labeled classes. Figure 4(b) shows how the Visual Prior Blending method can preserve the structure labeled classes in an image.
@@ 4.3.1 Effects of different methods:
The performance of the proposed methods is summarized in Table 3. When using the Baseline with generated prompts, the performance is only as low as 42.15% mIoU, which is lower than when training on the original data. This demonstrates the essential nature of the proposed methods when generating synthetic data. Our method yields a 50.27% mIoU, showing its effectiveness. Additionally, we experimented with combining the Post Filter with the Category-Calibrated CLIP Rank [Fang et al., 2024] for generating synthetic data. The result of this combination is higher, which shows that our method can combine filters from previous studies [Fang et al., 2024, Wu et al., 2024] to improve the performance.
@@ 4.3.2 Effect of number of synthetic data:
4.3.3 Text prompt seclection:
@@ 4.3.4 Effect of different trade-off scales:
Trade-off scales are utilized to combine the visual prior of the image with the semantic segmentation map presented in Section 3.2.2. We tested various scales and documented the results in Table 6. The outcomes indicate that the scale $\omega_{1}$=0.7 and $\omega_{2}$=0.9 yields the best results, with $\omega_{2}$ enabling proper localization of the labeled classes. On the other hand, the scale $\omega_{1}$=0.7 retains the general content of the image without needing to be as detailed as the original image.

### v4

@@ 1 INTRODUCTION
  • Visual Prior Blending: This method blends the visual priors from the original image and the segmentation masks. It aims to balance the information between the original image and the labeled objects, ensuring that the generated objects maintain the structure defined by the segmentation labels.
@@ 3 METHOD
Our proposed image data augmentation pipeline is shown in Figure 1, which consists of three main components: (1) Text prompt construction, (2) Visual Prior Blending, and (3) Controllable Diffusion Generation. Class-Prompt Appending append “class prompt” including the classes visible in the image with “caption” generated from the Image Captioning model. Visual Prior Blending method combines the visual pre-information of the real image and the segmentation map. The results of the above two methods are fed into the Controllable Stable Diffusion model to generate the synthetic image. In addition, to generate synthetic data from a given dataset, we use class balancing algorithm to generate data with even distribution among classes. Next, we provide a detailed description of each component.
@@ 3.2.2 Visual Prior Blending
In general, methods such as Line-Art Edge, Canny Edge, or HED soft edge all suffer from the limitation that the labeled classes in the image may be blurred or small in size, leading to inaccuracies in describing the structure of the labeled classes within the conditional image. Figure 4(a) shows an image produced using Line Art Detection. However, the edge results in this case are missing some details of the person, and the TV monitor is almost absent. The red shapes indicate the missing details in the image. This loss of information also occurs when using HED or Canny Edge. These weaknesses result in mislabeling in the synthetic image compared to the original image. We observed that although the segmentation labels of real images cannot fully describe an image’s content, they provide accurate information about the labeled classes. Based on this observation, we propose combining the real image’s prior visualization with the labels before feeding them into the controlled image generation model. The blending of the prior visual image $\mathcal{I}_{i}$ ($\mathcal{V}^{\mathcal{I}}_{i}$) and prior visual segmentation label $\mathcal{S}_{i}$ ($\mathcal{V}^{\mathcal{S}}_{i}$) ensures that the generated image has a clear layout and well preserves structures the class labeled information. Our proposed blends $\mathcal{V}^{\mathcal{I}}_{i}$ and $\mathcal{V}^{\mathcal{S}}_{i}$ by a weighted sum:
@@
With $\omega_{1}$, $\omega_{2}$ being the trade-off scales when combining $\mathcal{V}^{\mathcal{I}}_{i}$ and $\mathcal{V}^{\mathcal{S}}_{i}$. This blending results in a prior visual that is clear in content and complete information about labeled classes. Figure 4(b) shows how the Visual Prior Blending method can preserve the structure labeled classes in an image.
@@ 4.3.1 Effects of different methods:
The performance of the proposed methods is summarized in Table 3. When using the Baseline with generated prompts, the performance is only as low as 42.15% mIoU, which is lower than when training on the original data. This demonstrates the essential nature of the proposed methods when generating synthetic data. Our method yields a 50.27% mIoU, showing its effectiveness. Additionally, we experimented with combining the Post Filter with the Category-Calibrated CLIP Rank [Fang et al., 2024] for generating synthetic data. The result of this blending is higher, which shows that our method can combine filters from previous studies [Fang et al., 2024, Wu et al., 2024] to improve the performance.
@@ 4.3.2 Effect of number of synthetic data:
4.3.3 Text prompt selection:
@@ 4.3.4 Effect of different trade-off scales:
Trade-off scales are utilized to blend the visual prior of the image with the semantic segmentation map presented in Section 3.2.2. We tested various scales and documented the results in Table 6. The outcomes indicate that the scale $\omega_{1}$=0.7 and $\omega_{2}$=0.9 yields the best results, with $\omega_{2}$ enabling proper localization of the labeled classes. On the other hand, the scale $\omega_{1}$=0.7 retains the general content of the image without needing to be as detailed as the original image.
@@ 5.2 Conclusion
6 ACKNOWLEDGMENTS
This research is funded by University of Information Technology-Vietnam National University of Ho Chi Minh city under grant number D1-2024-76.